# Three-Field String Inflation with Perturbative Corrections: Dynamics and Implications

### v1

@@ 3.2 Multi-field perturbations
In case of multi-field inflation, it is standard to decompose the linear equation of motion for Mukhanov-Sasaki variable in kinematic basis i.e. along $(T^{a},N^{a},B^{a})$. The projection of field fluctuations $Q^{a}$ along the unit vectors are $Q_{i}=(Q_{T},Q_{N},Q_{B})\equiv(Q^{a}T_{a},Q^{a}N_{a},Q^{a}B_{a})$. The equations of motion in terms of $Q_{i}$ are Pinol:2020kvw ; Aragam:2023adu ; Aragam:2024nej :
@@ 4 Three-field analysis
In this section, we will thoroughly analyze the complete multi-field inflationary evolution and the detailed predictions of the model. To achieve this, we begin by outlining the consistency checkpoints that must be met in a viable string theory model building.
@@ 4.1 Cosmological evolution
$V$ is the scalar potential in (2.37) which consists of leading order $\alpha^{\prime 3}$ BBHL-term, log-loop corrections, higher derivative $F^{4}$ corrections and a constant uplift term resulting to a Minkowski-like vacuum of the potential at its global minimum. We are agnostic about the exact computation of the uplift term, it can either be a D-term uplift Burgess:2003ic , a T-brane uplift Cicoli:2015ylx or an anti-brane uplift Kachru:2002sk ; Crino:2020qwk ; Vafa:2005ui ; Cicoli:2024bxw . We now present a three-field inflationary model where all the three fields $(\varphi_{1},\varphi_{2},\varphi_{3})$ are displaced from their respective minimum. We get sufficient efolds as these fields complete their journey towards their minima and oscillate there to start the reheating process [^6] Note that, we do not intend to study the reheating dynamics in this article. Below, we discuss one benchmark model which uses the following set of parameters:
@@
Table 3: Model parameters and initial conditions for inflation labelled by superscript 00.
@@
Table 4: The values of different mass scales relevant to model building are considered. The Hubble parameter, derived from (3.2), remains nearly constant throughout the inflationary evolution, which is why a wiggle sign has been included.
@@
Figure 4: Field evolutions of the three canonical scalar fields. $\epsilon_{H}$ stays less than $1$ throughout the evolution but $\eta_{H}$ crosses $1$ at $N=60$. The reason for the same have been explained in the text.
@@
If we assess the initial conditions for the inflatons and their respective minima in table 3, we see that the initial path of the inflationary trajectory starts on the left side of $\langle\varphi_{1}\rangle$ and $\langle\varphi_{2}\rangle$. While both $(\varphi_{1},\varphi_{2})$ are displaced away from their minima, if we choose $\varphi_{3}$ close to its minimum, we obtain an interesting three-field attarctor where all three-fields move together to give sufficient amount of efolds. As soon as the fields settle at their minima, inflation ends. In Figure 4, although along $\varphi_{1}$ and $\varphi_{2}$ the displacement in fields $(\Delta\varphi_{1,2})$ remain relatively small in most of their evolution, they do exhibit some movement, as evident from the inset image. The potential along the $\varphi_{3}$ is the steepest among the rest, it dominates the field trajectory in the beginning when the $(\varphi_{1},\varphi_{2})$ move relatively slower. Hence, we may conclude that most of the slow-roll dynamics upto $N=55$ is mostly a cumulative effect of these fields with contributions from $\varphi_{3}$ being the dominant one. An interesting observation here is that, although Table 4 shows that $\varphi_{3}$ has the lowest mass at the global minimum, as the inflaton evolves from a shifted position, this hierarchy reverses, making $\varphi_{3}$ the steepest among all.
@@
The trajectory of $\varphi_{3}$ shows a piculiar behavior because if we notice and compare between the path $\varphi_{3}$ follows and its value of the minimum from table 3 then we can see that they are not in accordance. Roughly $N=55$ onwards, we notice vigorous oscillations in both $\epsilon_{H},\eta_{H}$ followed by a softening where $\eta_{H}$ returns to slow-roll after a brief period of slow-roll violation. We have $\epsilon_{H}<1$ throughout but its amplitude increases beyond $N=55$. The oscillation in $(\epsilon_{H},\eta_{H})$ is a manifestation of the oscillation in $\varphi_{3}$ direction around $N\sim 55$.
@@
We can understand this behavior by looking at the field trajectories as well as the steepness of the potential along $\varphi_{3}$. The reason is $\varphi_{3}$ falls in a ridge of the potential. We notice that along the inflationary trajectory where $(\varphi_{1},\varphi_{2})$ are far away from their minima, the true minimum for $\varphi_{3}$ disappears and the ridge becomes unavoidable. Although, notice that the ridge is not at the point where the potential diverges which is at $\varphi_{3}\equiv\sqrt{4/3}$. The ridge acts like a false vacuum for $\varphi_{3}$ and when it falls there it leads to such oscillations. Now, as soon as $(\varphi_{1},\varphi_{2})$ move close to their minima along the trajectory, $\varphi_{3}$ recovers its actual minimum and $\varphi_{3}$ restarts its journey towards the true vacuum. Interestingly, it appears that after the oscillation the fields switch their roles, $(\varphi_{1},\varphi_{2})$ move faster whereas $\varphi_{3}$ appears relatively constant. For the last few efolds, all the fields become active to roll towards the global minimum, thereby increasing both $\epsilon_{H}$ and $\eta_{H}$, signalling the end of inflation. These behavior can also be seen in the figure below.
Figure 5: Velocity of the fields.
@@
As expected, $\varphi_{3}$ being the steepest among them, its velocity dominates over the other and as soon as it falls off the ridge the velocity decreases but does not become zero. When $\varphi_{3}$ oscillates at the ridge, $(\varphi_{1},\varphi_{2})$ become dominant and move towards their minimum. In summary, this attarctor solution we have discovered, gives a quasi-single-field behavior in the first part of inflation and quasi-double-field for the second period of inflation. We will confirm this claim by looking at the projection of the masses of the scalar fields as well as perturbative masses of the modes throughout the inflationary evolution. We will discuss this in greater detail in the following subsections.
@@
Figure 6: Correct hierarchy of mass scales where $M_{s}>M_{KK}>m_{3/2}$. All the masses are expressed in Planck units.
@@
Figure 7: Hierarchy of mass scales
@@ 4.2 Cosmological parameters
$\xi^{\prime}=2a\Omega$ sources the coupling between curvature and entropic mode. $\Omega\sim 0$ naturally makes the coupling weak in our case. The masses of adiabatic and entropic perturbations can be derived from both (3.12) and the matrix (3.10). These quantities play a crucial role in determining the extent to which entropic perturbations influence inflationary perturbations, guiding us either deeper into a multi-field regime or away from it. The masses are defined as follows:
@@
We find that $\mathcal{M}_{\text{eff}}\sim M$, leading to $c_{s}\sim 1$ along the entire inflationary trajectory. Consequently, this results in negligible equilateral-type non-Gaussianity, with $f_{NL}\sim\mathcal{O}(10^{-8})$, as given by the relation in Aragam:2019omo :
@@
which holds in our case with $\mathcal{A}\sim\mathcal{O}(10^{-5})$. Therefore, given the small value of $\mathcal{A}$ as well as the mass of the adiabatic perturbation being the dominant one, we use the standard formulae for the cosmological parameters in terms of the slow-roll parameters (3.5),
@@
The amplitude of the power spectrum is given by,
@@
Figure 9: All the quantities are calculated at horizon crossing, which is roughly $50-60$ efolds before the end of inflation which is at $N_{\epsilon_{H}=1}=71.15$.
@@
Before ending this section, let us drop our final remark on the implications of the enhancement in the magnitude of turning rate and torsion as has been shown in figure 7. It is widely known in the literature Braglia:2020eai ; Ishikawa:2021xya ; Palma:2020ejf ; Parra:2024usv that models for which second slow-roll paramaters i.e. $\eta_{H}$ is larger than $\epsilon_{H}$, it can potentially gives rise to efficient production of primordial black holes. The field behavior reflected in our model reveals a similar behavior like double-inflation. In double inflation, the process occurs in two stages. Initially, one field rolls down to its minimum while the second field remains frozen. In the later stage, the frozen field begins to roll, ultimately reaching its minimum and bringing inflation to an end. Similarly, in the case of the three-field inflation discussed in this paper, we observe a quasi-double inflation behavior, where the first phase of inflation is mostly driven by one field, while the second phase is governed by the other two fields as we can observe in figure 4. Although, in this case, the first phase of slow-roll takes place along the steep direction of the potential.
Referring to figure 8, we observe that during horizon crossing $(11.15<N<21.15)$, the mass of the adiabatic mode, denoted as $\{M_{TT}^{2},m_{T}^{2}\}$, remains dominant. But on a smaller scale however, when the inflationary trajectory enters its second phase, characterized by two-field dynamics, where $(\varphi_{1},\varphi_{2})$ becomes more active compared to $\varphi_{3}$ – the masses of the entropic perturbations become dominant. This implies that the effects of turns and torsions must be accounted for, requiring a simultaneous analysis of both adiabatic and all entropic perturbations to determine the power spectrum of scalar and tensor modes. This highlights one of the most significant aspects of our paper.
Primordial black holes, considered potential progenitors of dark matter, have been widely studied in recent years, including a recent investigation within a three-field model in Aragam:2024nej where the authors presented a toy model. We emphasize that our work is among the first to explore a three-field inflation scenario with non-trivial features at small scales — motivated by type-IIB string theory flux compactification where our inflatons are four-cycles of a K3-fibred Calabi-Yau. The potential is a combination of leading order and one sub-leading terms in the inverse string expansion, and 1-loop correction terms. To ensure a nearly Minkowskian global minimum, a constant term has also been added. The fields are canonical and dimensionless in Planck units. We have done a thorough consistency checks to ensure the reliability of the 4d EFT. We plan to revisit the detailed study of primordial black holes as well as secondary induced gravitational waves in a forthcoming companion paper.
@@ 5 Conclusions
In this paper, we have presented a string motivated three-field inflationary construction for the first time where the hierarhcy between first and second slow-roll parameters become evident. The second slow-roll parameter dominates the first one, clearly hinting towards non-trivial dynamics at small scales – at the same time being consistent with Planck data at the horizon crossing. The scalar potential also deviates from the usual pLVS Antoniadis:2018hqy or fibre inflationary setup by their interplay between subleading corrections. The analysis is focused on the implications of the logarithmic loop corrections in the stabilization procedure, where the subleading $F^{4}$ higher derivative corrections were utilized to uplift the flat direction. We also added a small uplifting term to ensure the presence of a Minkowski vacuum.
@@
The above method is compatible with the large volume scenario (LVS) in which the compatification volume is stabilized to large values of order $>\mathcal{O}(10^{6})$ in accordance with the semi-classical approximation and trust the underlying 4D EFT. The model reflects the correct hierarchy of mass scales, i.e. $M_{pl}>M_{s}>M_{KK}>M_{inf}$ in all the benchmark models displayed. As far as the inflationary setup is concerned, we present a multi-field inflation model where all the scalar fields are active during inflation. The inflation takes place along $(\varphi_{1},\varphi_{2},\varphi_{3})$ directions which span the canonical basis constructed out of several four-cycle volume of the fibre modulus $\tau_{7}$, the volume of the base $\tau_{6}$ and the blow-up modulus $\tau_{1}$. The three-dimensional scalar potential, defined in a canonical basis within the scalar field space, is sufficiently flat to sustain the necessary number of e-folds. Inflation starts at a position where all the scalars are shifted from their global minimum. However, it features a ridge-valley along the $\varphi_{3}$ direction when the other fields are away from their minimum. The potential along that direction becomes steeper than the other two directions $(\varphi_{1},\varphi_{2})$.
@@
The minimum along $\varphi_{3}$ vanishes at $\{\varphi_{1}^{\text{shifted}},\varphi_{2}^{\text{shifted}}\}$ forcing $\varphi_{3}$ to fall into the ridge and oscillate there – acting like a false vacuum for $\varphi_{3}$. In the second stage, as $\varphi_{3}$ stops its oscillation, the other two fields starts its journey with more kinetic energy towards their minima. As the value of $(\varphi_{1},\varphi_{2})$ start to get near their minima, the true minimum for $\varphi_{3}$ is restored and $\varphi_{3}$ slowly transitions there as we can see in figure 5. This oscillation exhibits a non-trivial feature to which we now turn our attention. Due to the flat metric, as expected, throughout the slow-roll evolution, the model exhibits minor deviation from a geodesic trajectory. However, when $\varphi_{3}$ oscillates and the other two fields start to roll, it leads to high value of turing rate. The torsion which takes care of how much non-planar behavior is present in the trajectory, despite being small for most of the evolution, increases right after the onset of oscillation.
@@
We can notice the same even at the level of the masses of the adiabatic and entropic modes— they switch their roles as can be seen in figure 8. The domination of entropic mass at small scales indicates the fact that at the level of perturbation, three field analysis will become unavoidable. This type of features are closely related to the production of primordial black holes at small scales – at the time of radiation domination. Besides, it can also lead to non-trivial footprint on the scalar induced gravitational waves. We plan to return to this detailed study in a companion paper. We emphasize that this trajectory is very generic for this class of models, depending on the initial condition one can easily achieve this three-field attractor.
@@
The inflationary model presented in this paper is consistent with the data, however it has a few areas for further refinement. The value of the blow-up divisor at its minimum is large compared to the overall volume. However, we have made sure that the model still resides in the large volume regime. Note that, this is not the case for the dS models presented in table 1-2. The inflationary model is chosen so that the amplitude of the power spectrum is obtained without any fine-tuning or artificial scaling. Besides, the Kaluza-Klein and the string mass scales are almost comparable to each other in all the models— pushing us to the boundary of parametric control.
Many challenges remain as open questions with respect to the dynamics of this class of inflationary models. The exact computation of the values of $\chi$ and $\lambda$ is needed to unravel the robustness of the inflationary trajectories against additional string loop effects Gao:2022uop . An advantage of the proposed model is that it manages to have the EFT under control, despite the fact that it lacks an explicit mechanism of uplifting the vacuum to dS space. It would be worth studying the effects of the $\overline{D3}$ uplifting method Cicoli:2024bxw in the presence of the logarithmic loop corrections. We save these questions for a future work.

### v2

@@ 3.2 Multi-field perturbations
In case of multi-field inflation, it is standard to decompose the linear equation of motion for Mukhanov-Sasaki variable in kinematic basis i.e. along $(T^{a},N^{a},B^{a})$. The projection of field fluctuations $Q^{a}$ along the unit vectors are $Q_{i}=(Q_{T},Q_{N},Q_{B})\equiv(Q^{a}T_{a},Q^{a}N_{a},Q^{a}B_{a})$. The equations of motion in terms of $Q_{i}$ are Pinol:2020kvw ; Aragam:2023adu ; Aragam:2024nej ; Ashoorioon:2025iid :
@@ 4 Three-field analysis
In this section, we shall thoroughly analyze the complete multi-field inflationary evolution and the detailed predictions of the model. To achieve this, we begin by outlining the consistency checkpoints that must be met in a viable string theory model building.
@@ 4.1 Cosmological evolution
$V$ is the scalar potential in (2.38) which consists of leading order $\alpha^{\prime 3}$ BBHL-term, log-loop corrections, higher derivative $F^{4}$ corrections and a constant uplift term resulting to a Minkowski-like vacuum of the potential at its global minimum. We are agnostic about the exact computation of the uplift term, it can either be a D-term uplift Burgess:2003ic , a T-brane uplift Cicoli:2015ylx or an anti-brane uplift Kachru:2002sk ; Crino:2020qwk ; Vafa:2005ui ; Cicoli:2024bxw .
We now present a three-field inflationary model where all the three fields $(\varphi_{1},\varphi_{2},\varphi_{3})$ are displaced from their respective minimum. We get sufficient efolds as these fields complete their journey towards their minima and oscillate there to start the reheating process [^6] Note that, we do not intend to study the reheating dynamics in this article. Below, we discuss one benchmark model which uses the following set of parameters:
@@
Table 3: The table contains the model parameters and initial conditions. The latter is labelled by superscript 00.
@@
Table 4: Mass scales follow the correct hierarchy. The Hubble parameter, derived from (3.2), remains nearly constant throughout the inflationary evolution, which is why a wiggle sign has been included.
@@
Figure 4: Field evolutions of the three canonical scalar fields. $\epsilon_{H}$ stays less than $1$ throughout the evolution but $\eta_{H}$ crosses $1$ at $N=55$. The reason for the same have been explained in the text.
@@
If we assess the initial conditions for the inflatons and their respective minima in table 3, we see that the initial path of the inflationary trajectory starts on the left side of $\langle\varphi_{1}\rangle$ and $\langle\varphi_{2}\rangle$. While both $(\varphi_{1},\varphi_{2})$ are displaced away from their minima, if we choose $\varphi_{3}$ close to its minimum, we obtain an interesting three-field attarctor where all three-fields move together to give sufficient amount of efolds. As soon as the fields settle at their minima, inflation ends. In Figure 4, although along $\varphi_{1}$ and $\varphi_{2}$ the displacement in fields $(\Delta\varphi_{1,2})$ remain relatively small in most of their evolution, they do exhibit some movement, as evident from the inset images. The potential along the $\varphi_{3}$ is the steepest among the rest, it dominates the field trajectory in the beginning when the $(\varphi_{1},\varphi_{2})$ move relatively slower. Hence, we may conclude that most of the slow-roll dynamics upto $N=55$ is mostly a cumulative effect of these fields with contributions from $\varphi_{3}$ being the dominant one. An interesting observation here is that, although Table 4 shows that $\varphi_{3}$ has the lowest mass at the global minimum, as the inflaton evolves from a shifted position, this hierarchy reverses, making $\varphi_{3}$ direction the steepest among all.
@@
The trajectory of $\varphi_{3}$ shows a piculiar behavior because if we notice and compare between the path $\varphi_{3}$ follows and its value of the minimum from table 3 then we can see that the field is not directly going towards its minimum. Roughly $N=55$ onwards, we notice vigorous oscillations in both $\epsilon_{H},\eta_{H}$ followed by a softening where $\eta_{H}$ returns to slow-roll after a brief period of slow-roll violation. We have $\epsilon_{H}<1$ throughout but its amplitude increases beyond $N=55$. The oscillation in $(\epsilon_{H},\eta_{H})$ is a manifestation of the oscillation in $\varphi_{3}$ direction around $N\sim 55$.
@@
We can understand this behavior by looking at the field trajectories as well as the steepness of the potential along $\varphi_{3}$. The reason is $\varphi_{3}$ falls in a ridge of the potential. We notice that along the inflationary trajectory where $(\varphi_{1},\varphi_{2})$ are far away from their minima, the true minimum for $\varphi_{3}$ disappears and the ridge becomes unavoidable. Although, notice that the ridge is not at the point where the potential diverges which is at $\varphi_{3}\equiv\sqrt{4/3}$. The ridge acts like a false vacuum for $\varphi_{3}$ and when it falls there it leads to such oscillations [^7] Note that, it may be possible that this oscillation along $\varphi_{3}$ may trigger an early preheating. However, since slow-roll resumes after the end of oscillation, leading to the continued expansion of spacetime, any particles generated during this period will ultimately be diluted away.. Now, as soon as $(\varphi_{1},\varphi_{2})$ move close to their minima along the trajectory, $\varphi_{3}$ recovers its actual minimum and $\varphi_{3}$ restarts its journey towards the true vacuum. Interestingly, it appears that after the oscillation the fields switch their roles, $(\varphi_{1},\varphi_{2})$ move faster whereas $\varphi_{3}$ appears relatively slower, see figure 5. For the last few efolds, all the fields become active to roll towards the global minimum, thereby increasing both $\epsilon_{H}$ and $\eta_{H}$, signalling the end of inflation.
Figure 5: Speed of the inflatons.
@@
As expected, $\varphi_{3}$ being the steepest among them, its velocity dominates over the other and as soon as it falls off the ridge the velocity decreases but does not become zero. When $\varphi_{3}$ oscillates at the ridge, $(\varphi_{1},\varphi_{2})$ become dominant and move towards their minimum. In summary, this attarctor solution we have discovered, gives a quasi-single-field behavior in the first part of inflation and quasi-double-field for the second period of inflation. We shall confirm this claim by looking at the projection of the masses of the scalar fields as well as perturbative masses of the modes throughout the inflationary evolution. We shall discuss this in greater detail in the following subsections.
@@
Figure 6: Correct hierarchy of mass scales with $M_{s}>M_{KK}>m_{3/2}$. All the masses are expressed in Planck units.
@@
Figure 7: The dimensionless turning rate and torsion
@@ 4.2 Cosmological parameters
$\xi^{\prime}=2a\Omega$ sources the coupling between curvature and entropic mode. $\Omega\ll 1$ naturally makes the coupling weak in our case. The masses of adiabatic and entropic perturbations can be derived from both (3.12) and the matrix (3.10). These quantities play a crucial role in determining the extent to which entropic perturbations influence the whole inflationary perturbations, guiding us either deeper into a multi-field regime or away from it. The masses are defined as follows:
@@
We find that $\mathcal{M}_{\text{eff}}\sim M$, leading to $c_{s}\sim 1$ along the entire inflationary trajectory. Consequently, this results in negligible equilateral-type non-Gaussianity, with $f_{NL}\sim\mathcal{O}(10^{-8})$, as given by the relation Aragam:2019omo :
@@
which holds true in our case with $\mathcal{A}\sim\mathcal{O}(10^{-5})$. Therefore, given the small value of $\mathcal{A}$ as well as the mass of the adiabatic perturbation being the dominant one, we use the standard formulae for the spectral index and spectral tilt in terms of the slow-roll parameters (3.5):
@@
The amplitude of the scalar power spectrum $(\Delta_{s}^{2})$ and tensor power spectrum $(\Delta_{t}^{2})$ are given by:
@@
Figure 9: Scalar power spectrum $(\Delta_{s}^{2})$, spectral index $(n_{s})$ and spectral tilt $(r)$ are calculated at horizon crossing, that is roughly $50-60$ efolds before the end of inflation which is at $N_{\epsilon_{H}=1}=71.15$.
@@
Before ending this section, let us drop our final remark on the implications of the enhancement in the magnitude of turning rate and torsion as has been shown in figure 7. It is widely known in the literature Braglia:2020eai ; Palma:2020ejf ; Fumagalli:2020adf ; Fumagalli:2020nvq ; Ishikawa:2021xya ; Fumagalli:2021mpc ; Parra:2024usv that models for which second slow-roll paramaters i.e. $\eta_{H}$ is larger than $\epsilon_{H}$, together with a brief period of slow-roll violation is present – the cumulative effect can potentially give rise to efficient production of primordial black holes. The field behavior reflected in our model reveals a similar behavior like double-inflation. In double inflation, the process occurs in two stages. Initially, one field rolls down to its minimum while the second field remains frozen. In the later stage, the frozen field begins to roll, ultimately reaching its minimum and bringing inflation to an end. Similarly, in the case of the three-field inflation discussed in this paper, we observe a quasi-double inflation, where the first phase of inflation is mostly driven by one field, while the second phase is governed mostly by the other two fields as we can observe in figure 5. First phase of slow-roll takes place along the steep direction of the potential.
Referring to figure 8, we observe that during horizon crossing $(11.15<N<21.15)$, the mass of the adiabatic mode, denoted as $\{M_{TT}^{2},m_{T}^{2}\}$, remains dominant. But on a smaller scale however, when the inflationary trajectory enters its second stage, characterized by two-field dynamics, where $(\varphi_{1},\varphi_{2})$ becomes more active compared to $\varphi_{3}$ – the masses of the entropic perturbations become dominant. This implies that the effects of turns and torsions must be accounted for, requiring a simultaneous analysis of both adiabatic and all entropic perturbations to determine the power spectrum of scalar and tensor modes.
Primordial black holes, considered potential progenitors of dark matter, have been widely studied in recent years, including a recent investigation within a three-field model in Aragam:2024nej where the authors presented a toy model. We emphasize that our work is among the first to explore a three-field inflation scenario with non-trivial features at small scales — motivated by type-IIB string theory flux compactification where our inflatons are four-cycles of a K3-fibred Calabi-Yau. The potential is a combination of leading order BBHL term, higher derivative $F^{4}$ term, and 1-loop correction terms. To ensure a nearly Minkowskian global minimum, a constant term has also been added. The fields are canonical and dimensionless in Planck units. We have performed a thorough consistency check to ensure the reliability of the 4d EFT. We plan to revisit the detailed study of primordial black holes as well as secondary induced gravitational waves in a forthcoming companion paper.
@@ 5 Conclusions
In this paper, we have presented a three-field inflationary construction in string theory flux compactification for the first time where the hierarhcy between first and second slow-roll parameters with a transient violation of slow-roll becomes evident. The model is consistent with the Planck data at the horizon crossing. The scalar potential deviates from the usual pLVS Antoniadis:2018hqy or fibre inflationary setup in their interplay of subleading corrections. The analysis is focused on the implications of logarithmic loop corrections and higher derivative $F^{4}$ correction in the moduli stabilization procedure. We also added a small uplifting term to ensure the presence of a Minkowski vacuum.
@@
The volume of the CY is stabilized to a large value of order $\mathcal{V}>\mathcal{O}(10^{6})$ (see table 1) and hence we can trust the underlying 4D low energy EFT in making the correction terms subleading in $\alpha^{\prime}$ and $g_{s}$ expansions. The model reflects the correct hierarchy of mass scales, i.e. $M_{pl}>M_{s}>M_{KK}>M_{inf}$ in all the benchmark models displayed in table 1. As far as the inflationary setup is concerned, we present a multi-field inflation model where all the scalar fields are active during inflation. Inflation occurs along $(\varphi_{1},\varphi_{2},\varphi_{3})$ directions which span the canonical field space metric, constructed out of several four-cycles such as fibre modulus $\tau_{7}$, the volume of the base $\tau_{6}$ and the blow-up modulus $\tau_{1}$. Inflation starts at a position where all the scalars are shifted from their global minimum. However, it features a ridge-valley along the $\varphi_{3}$ direction when the other fields are away from their minimum. The potential in that direction becomes steeper than in the other two directions $(\varphi_{1},\varphi_{2})$.
After analyzing the field trajectory, we confirm that inflation along this attractor occurs in two distinct stages. The first stage is mostly dominated by $\varphi_{3}$ because the potential in this direction is the steepest. The minimum along $\varphi_{3}$ disappears at $\{\varphi_{1}^{\text{shifted}},\varphi_{2}^{\text{shifted}}\}$, but forms a ridge —compelling $\varphi_{3}$ to settle into it and oscillate, effectively behaving as a false vacuum for $\varphi_{3}$. The ridge continues to stay until the other two fields start to roll near their respective minima — signalling the onset of the second stage of inflation. When comparing the speed of the inflatons, the inflationary trajectory in the first stage is largely dominated by $\varphi_{3}$— leading to a behavior of quasi-single field. The horizon crossing of the pivot scale $(k_{\star})$ occurs in the first stage and the model is observationally viable.
In the second stage, as $\varphi_{3}$ stops its oscillation, the other two fields start their journey towards their minima. As soon as $(\varphi_{1},\varphi_{2})$ start to get near their minima, the true minimum for $\varphi_{3}$ is restored and all the fields slowly roll towards their respective minimum we can see in figure 5. Hence, in the second stage, the field velocity is mostly dominated by $(\varphi_{1},\varphi_{2})$. We can that these two stages exhibit a quasi-double-inflation like behavior where the first stage is dominated by $\varphi_{3}$ and the second by $(\varphi_{1},\varphi_{2})$.
@@
This oscillation demonstrate a non-trivial feature to which we now turn our attention. Due to the flat metric, as expected, during the first stage, the model exhibits minor deviation from a geodesic trajectory. However, when $\varphi_{3}$ oscillates and the other two fields start to roll, it leads to high value of turning rate. The torsion which takes care of how much non-planar behavior is present in the trajectory, despite being small for most of the evolution, increases right after the onset of oscillation.
@@
We can notice the same even at the level of the masses of the adiabatic and entropic modes— they switch their roles as can be seen in figure 8. The domination of entropic mass at small scales indicates the fact that at the level of perturbation, three field analysis will become unavoidable. This type of features are closely related to the production of primordial black holes at small scales – possibly at the time of radiation domination. Besides, it can also lead to non-trivial footprint on the scalar induced gravitational waves. We plan to return to this detailed study in a companion paper. We emphasize that this trajectory is very generic for this class of models, depending on the initial condition one can easily achieve this three-field attractor.
@@
There exists few areas in the inflationary model and the examples presented in table 1-2 that call for further refinement. The example in table 3, the vev of the blow-up modulus is calculated to be of order $\mathcal{O}(10^{6})$ whereas the overall volume is $\mathcal{O}(10^{5})$. However, we have made sure that the model still resides in the large volume regime throughout the inflationary evolution. Note that, this is not the case for the dS models presented in table 1-2. The inflationary model is chosen so that the amplitude of the power spectrum is obtained without any fine-tuning or artificial scaling. Besides, the Kaluza-Klein and the string mass scales are almost comparable to each other in all the models— pushing us to the boundary of parametric control. In all the examples de Sitter vacua presented in this paper, large value of blow-up modulus invariably implies a vanishing non-perturbative effects. Since, these effects do not play an active role in moduli stabilization, we do not face any issue.
The log-loop corrections are presented in a conjectured form for a K3-fibred CY, based on calculations performed in the case of a torus. In addition, since the stabilization proceeds via loop corrections, the axionic moduli remains massless because the Kähler potential retains the axionic shift symmetry. Subsequently, these axions can play non-trivial roles in driving several post-inflationary aspects such as reheating, and can source dark radiation and dark matter (see Cicoli:2012aq ; Cicoli:2023opf for details). Moreover, it has been emphasized in Blanco-Pillado:2009dmu that if one considers the inflationary evolution of the axions, it usually gives rise to a greater variety of trajectories — with cases where the axionic evolution increase the effective number of efolds. However, addressing the stabilization, evolution and dynamics of the axions are beyond the scope of this paper. Furthermore, exact computation of the values of $\chi$ and $\lambda$ is needed to unravel the robustness of the inflationary trajectories against additional string loop effects Gao:2022uop . We save these questions for a future work.